# A sixteen multiple-amplifier-sensing CCD and characterization techniques targeting the next generation of astronomical instruments

### v1

@@ 4.3 Average GRE/GRI of the amplifier chain
4.4 Non-constant GRE/GRI
It is well known that CTI/CTE has a dependence on the charge packet size[16] and also on other factors such as temperature [17]. It is expected that the same nonlinear behavior could be present in GRE/GRI measurements. In this scenario, the recursive Eq.(10) is still valid, but the system becomes a time-variant system; therefore, the Z-transform analysis is not valid. Nevertheless, an algorithmic approach using the recursion could be implemented. Appendix B shows the recursive algorithm implemented to compare with our analytical model.
@@ 5 Experimental results
The sensor under study possesses $N_{a}=16$ non-destructive readout stages on the same serial line. A microscope image of the output stage of the MAS-CCD is shown in Fig. 4a. The figure displays the direction of charge movement through the multiple inline amplifiers. There is a fixed pixel separation between each output stage. The active region is located in the bottom left corner, with the serial register indicated. The sensor consists of a 1024 $\times$ 512 array of 15 $\mu$m $\times$ 15 $\mu$m pixels and a substrate thickness of 675 $\mu$m. The output amplifiers are separated by $15$ pixels. The pixels in the prescan are used to bend the serial register by 90 degrees as shown in the picture. The sensor is operated at 140K using a polycold cryocooler and in fully depletion mode using a substrate voltage of 70V.
@@ 5.1 Readout noise studies
This section demonstrates the noise reduction properties of the sensor operated in different conditions of integration time and the number of samples. Figure 5(a) shows the readout noise performance of the 16 amplifiers as a function of the readout integration time of the DSI operation using $N_{s}$ = 1 per amplifier. Similar operation is obtained from all of them. At the same time, the plot shows the measured noise after combining the information from all the channels with equal weights, following Eq.(1). The plot also shows the expected noise performance after this calculation which is given by Eq.(2). Both quantities have a good agreement.
Figure 5: a) Noise performance for different Integration times $T_{s}$ of all the channels, and combination and expected performance. b) Noise reduction of all the channels as a function of the Skipping samples $N_{s}$ and the standard and expected combination of the channels, for a integration time $T_{s}=16.66\mu$s.
The noise performance was also measured as a function of the number of samples per amplifier ($N_{s}$) using an integration time of $T_{s}$ = 16.6 $\mu$s. Figure 5(b) summarizes these results. All the amplifiers show similar noise performance. The plot also shows in circled markers the resultant noise when the information from the channels is averaged using Eq. (1). The combined measurement shows good agreement with the expected performance (square markers) from Eq. (2). The expected combined noise value for Ns=1 is used to extrapolate to larger Ns dividing by $\sqrt{N_{s}}$ giving a straight line in the plot, which is indicated as a green line in Fig.5(b). The line reveals that the measured performance follows the expected reduction fraction.
@@ 5.2 Performance at targeted speeds and implementation of processing techniques
Another way to read the sensor is to allow each amplifier to take several samples. The green cross markers in Fig. 6 show the noise performance of the system when multiple samples are taken on each amplifier. To compute the final pixel value, the first step is to average the samples from individual amplifiers using weights of 1/$N_{s}$. Secondly, the resulting values from the amplifiers are optimally combined following the recipes in Sec. 3.1 and 3.2. These measurements were performed by increasing the number of samples per amplifier from the original readout speed of $\sim$72k pixels/sec, indicated in the plot with the label $N_{s}=1$. The dashed lines of the same color visually connect these points in the plot. It is interesting to note that increasing $N_{s}$ gives a better noise performance than the red square dots, which represent images processed with the same techniques. It is expected that the performance of the multiple skipping samples for the same total pixel time could give a different noise performance due to the different effective frequency responses. Please see the appendix A for a discussion about the expected frequency of the system under different parameter conditions. One interesting feature to note is that the measured noise performance from multiple samples follows the expected reduction by the factor $\sqrt{N_{s}}$ from the starting point at $N_{s}=1$. Applying the noise decorrelation technique (discussed in Sec. 3.3) on the multiple $N_{s}$ measurements we obtained the performance shown with plus markers in light blue. This processing has little effect on the performance indicating that there is little contribution of common noise to subtract between amplifiers. The multiple sampling operation per amplifier provides an extra degree of freedom for the optimization of the readout operation of this kind of sensor. If the multiple sampling readout of the entire pixel is a limitation for the measurements of all the pixels in the active regions, a selective strategy with regions of interest as in [22] could be used to increase the signal-to-noise ratio of interesting pixels.
@@
5.3 Gate removal efficiency
@@
As discussed before, the GRE has a strong dependency on the operation of the sensor: voltages, size of the charge packet, and the speed of the clocks. The goal of this section is to study experimentally the GRE for different cases of $\epsilon$ and provide ideas on how to operate this sensor compared to the regular single-stage detectors. For this task, we show three witness cases which are represented by the plots in Fig. 8. The plots show the measured pixel value as a function of the column index. Each trace is the average of 15 rows that are averaged to reduce the uncertainty in the pixel measurements. The plots show the transition from the active to the overscan pixels for all the channels of the sensor, producing a negative step of charge. The measured pixels are shown in circle markers and a dashed line with the same color is used to connect them and to help the reader to follow the abrupt transition between the regions. The sensor was illuminated prior to the readout and areas with uniform light exposure were used. A moderate amount of charge was intentionally used in the pixels to show the behavior in a realistic scenario. The black dots in some of the plots represent the fit of our model to the data. Each plot has a title identifying the most important conditions of voltages in the gates of the sense node (Vref) and the low levels of the PS (VPS) and H (VH) clocks, as well as the amount of charge used in this case. To generate the plots, the offset of each image is subtracted and the values of the pixels are normalized by the amount of charge in the active pixels a few pixels before the transition to the overscan. This representation helps to easily infer the $\epsilon$ values from the plots using the normalized value of the pixels in the overscan. The following scenarios were considered:
@@
  • GRI below the precision of the method. In Fig. 8a, the first pixel in the overscan (i.e., the first pixel after the abrupt transition) does not have an evident trace of charge for all the channels. The inlet shows a zoom-in in the overscan regions. The first pixel seems to have a statistic similar to the rest of the pixels of the overscan. In this case, $\epsilon<5\times 10^{-5}$.
@@
  • Moderate GRI. Figure 8b shows the case where the first pixel in the overscan starts growing as a function of the number of amplifiers proportionally, such as in Eq.(17). In this case, $\epsilon=5\times 10^{-3}$ from the fitted model, which is shown with the black dots.
@@
  • Large GRI. Figure 8c shows an extreme case where the PS and VH low-level voltages of the clocks are set close to the gate voltage of the sense node to force a bad charge removal from the sense node (very large GRI). The intermediate pixels in the transitions from the active to the overscan regions are due to this poor removal capability. In this case, the GRI is very severe, and the pixels in the overscan do not show a proportional growth with the amplifier number. In this extreme case, the best-fit model was obtained by assuming a $\epsilon$ with proportional dependence $\epsilon=2.26\times 10^{-6}Q+3\times 10^{-4}$. The algorithmic model, presented in Appendix B, was used for the calculation.
@@
Figure 8: Measured charge value as a function of the image’s column index for the sixteen MAS-CCD channels. In particular, the plot shows the transition between the pixel in the active region and the overscan region. The different plots show different levels of GRE effect on the measured signal.
@@
Another important aspect to note here is that the removal of the charge from the sense nodes requires voltages below the sense node reference voltage. For the case of CCD collecting holes, this means that the voltages should be more negative than the sense node gate voltage. The current limitation of our hardware does not allow us to explore clock voltages below -10 volts; there is an opportunity for future measurements to improve the GRE by using smaller clock values. Another aspect to incorporate in future instrumentation of this kind of sensor is the use of a clock to set voltages in the sense node instead of a fixed bias voltage (as currently being used). A clock would allow an increase in the voltage in the sense node when transferring the change to increase the gradient towards the PS gate.
@@ Acknowledgments
The fully depleted Skipper-CCD was developed at Lawrence Berkeley National Laboratory, as were the designs described in this work. The CCD development work was supported in part by the Director, Office of Science, of the U.S. Department of Energy under No. DE-AC02-05CH11231. This research has been sponsored by the Laboratory Directed Research and Development Program of Fermi National Accelerator Laboratory (L2019-011, L2022.053), managed by Fermi Research Alliance, LLC for the U.S. Department of Energy. This research has been partially supported by Guillermo Fernandez Moroni’s DOE Early Career research program.
@@ Appendix A Impulse and frequency response of the MAS-CCD to common noise
The Dual-Slope Integrator (DSI) impulse response and frequency response are [13, 11]
$h_{DSI}\left(t\right)=\begin{cases}\dfrac{A}{T_{s}},\hskip 5.69046pt0\leq t\leq 0+T_{s}\\ -\dfrac{A}{T_{s}},0+T_{s}\leq t\leq 0+2T_{s}\\ 0,\hskip 28.45274pt\text{in other case}\end{cases}$ (19)
@@
where $T_{s}$ is the integration time for one sample and $t_{0}$ is an arbitrary start point of the integration.
@@
$H_{SKP}\left(f\right)=\dfrac{H_{DSI}\left(f\right)}{N_{s}}\left|\frac{\sin\left(2\pi NT_{s}f\right)}{\sin\left(2\pi T_{s}f\right)}\right|$
@@
h_MAS( t) =1Na1Ns∑^N_a-1_m=0∑^N_s-1_n=0h_DSI( t-mΔt -2nT_s)
where $\Delta t$ model the required transfer time and intermediate readout time in between stages.
@@
Appendix B GRE recursive algorithm
GRE recursive algorithm that was used to compare with the analytical calculations in Sec. 4. The algorithm also provides a way to implement GRI values that are dependent of the charge in the sense node.
Listing 1: GRI implementation in python.
@@ Appendix B GRE recursive algorithm
def pixelsAfterNStages(pix, Na):
@@
measPix = np.zeros((Na, Npix))
@@
measPix[j, n] = inPix[n]
# Gate Removal Inefficiency calculation (GRI). epsilon is
# the GRI value as used in the text of the article.

### v2

@@ 4.3 Average NRE/NRI of the amplifier chain
4.4 Non-constant NRE/NRI
It is well known that CTI/CTE has a dependence on the charge packet size[20] and also on other factors such as temperature [21]. It is expected that the same nonlinear behavior could be present in NRE/NRI measurements. In this scenario, the recursive Eq.(10) is still valid, but the Z-transform analysis cannot be used. Nevertheless, an algorithmic approach using the recursion could be implemented. Appendix B shows a complete and fully functional example of the recursive algorithm implemented for the nonlinear case.
@@ 5 Experimental results
The sensor under study possesses $N_{a}=16$ non-destructive readout stages on the same serial line and operates by collecting holes. A microscope image of the output stage of the MAS-CCD is shown in Fig. 4a. The figure displays the direction of charge movement through the multiple inline amplifiers. The active region is located in the bottom left corner, with the serial register indicated. The sensor consists of a 1024 $\times$ 512 array of 15 $\mu$m $\times$ 15 $\mu$m pixels and a substrate thickness of 675 $\mu$m. The output amplifiers are separated by $15$ pixels. The pixels in the prescan are used to bend the serial register by 90 degrees as shown in the picture. The sensor is operated at 140K using a polycold cryocooler and in fully depletion mode using a substrate voltage of 70V.
@@ 5.1 Readout noise studies
This section demonstrates the noise reduction properties of the sensor operated in different conditions of integration time and the number of samples $N_{s}$ taken by each amplifier. Figure 5(a) shows the readout noise performance of the 16 amplifiers as a function of the readout integration time ($T_{s}$) of the DSI operation (mathematical model presented in Appendix A), using $N_{s}$ = 1 per amplifier. Similar operation is obtained from all of them. At the same time, the plot shows the measured noise after combining the information from all the channels with equal weights, following Eq.(1). The plot also shows the expected noise performance after this calculation which is given by Eq.(2), both quantities have a good agreement.
Figure 5: a) Noise performance for different Integration times $T_{s}$ of all the channels, and combination and expected performance. b) Noise reduction of all the channels as a function of the Skipping samples $N_{s}$ and the standard and expected combination of the channels, for an integration time $T_{s}=16.66\mu$s.
The noise performance was also measured as a function of the number of samples per amplifier ($N_{s}$) using an integration time of $T_{s}$ = 16.6 $\mu$s is shown in Fig. 5(b). All the amplifiers show similar noise performance. The plot also shows in circled markers the resultant noise when the information from the channels is averaged using Eq. (1). The combined measurement shows good agreement with the expected performance (square markers) from Eq. (2). The expected combined noise value for $N_{s}=1$ is used to extrapolate to larger $N_{s}$ dividing by $\sqrt{N_{s}}$ giving a straight line in the plot, which is indicated as a green line in Fig.5(b). The line reveals that the measured performance follows the expected reduction fraction.
@@ 5.2 Performance at targeted speeds and implementation of processing techniques
Another way to read the sensor is to allow each amplifier to take several samples. The light blue plus markers in Fig. 6 show the noise performance of the system when multiple samples are taken on each amplifier. To compute the final pixel value, the first step is to average the samples from individual amplifiers using weights of 1/$N_{s}$. Secondly, the resulting values from the amplifiers are optimally combined following the recipes in Sec. 3.1 and 3.2. These measurements were performed by increasing the number of samples per amplifier from the original readout speed of $\sim$72k pixels/sec, indicated in the plot with the label $N_{s}=1$. The dashed lines of the same color visually connect these points in the plot. It is interesting to note that increasing $N_{s}$ gives a better noise performance than the red square dots, which represent images processed with the same techniques. It is expected that the performance of the multiple skipping samples for the same total pixel time could give a different noise performance due to the different effective frequency responses. Please see the appendix A for a discussion about the expected frequency of the system under different parameter conditions. One interesting feature to note is that the measured noise performance from multiple samples follows the expected reduction by the factor $\sqrt{N_{s}}$ from the starting point at $N_{s}=1$.
Applying the noise decorrelation technique (discussed in Sec. 3.3) on the multiple $N_{s}$ measurements we obtained the performance shown with cross markers in green. This processing has little effect on the performance indicating that there is little contribution of common noise to subtract between amplifiers. The multiple sampling operation per amplifier provides an extra degree of freedom for the optimization of the readout operation of this kind of sensor. If the multiple sampling readout of the entire pixel is a limitation for the measurements of all the pixels in the active regions, a selective strategy with regions of interest as in [26] could be used to increase the signal-to-noise ratio of interesting pixels.
@@
5.3 Node removal efficiency
@@
As discussed in Sec. 4, the NRE has a strong dependency on the operation of the sensor: voltages, size of the charge packet, and the speed of the clocks. The goal of this section is to study experimentally the NRE for different cases of $\epsilon$ and provide ideas on how to operate this sensor compared to the regular single-stage detectors. For this task, we show three witness cases which are represented by the plots in Fig. 8. The plots show the measured pixel value as a function of the column index. Each trace is the average of 15 rows that are averaged to reduce the uncertainty in the pixel measurements. The plots show the transition from the active to the overscan pixels for all the channels of the sensor, producing a negative step of charge. The measured pixels are shown in circle markers and a dashed line with the same color is used to connect them and to help the reader to follow the abrupt transition between the regions. The sensor was illuminated prior to the readout and areas with uniform light exposure were used. A moderate amount of charge was intentionally used in the pixels to show the behavior in a realistic scenario. The black dots in some of the plots represent the fit of the proposed NRI model to the data. Each plot has a title identifying the most important conditions of voltages in the gates of the sense node (Vref) and the low levels of the PS (VPS) and H (VH) clocks, as well as the amount of charge used in each case. To generate the plots, the offset of each image is subtracted and the values of the pixels are normalized by the amount of charge in the active pixels a few pixels before the transition to the overscan. This representation helps to easily infer the $\epsilon$ values from the plots using the normalized value of the pixels in the overscan. The following scenarios were considered:
@@
  • NRI below the precision of the method. In Fig. 8a, the first pixel in the overscan (i.e., the first pixel after the abrupt transition) does not have an evident trace of charge for all the channels. The inlet shows a zoom-in in the overscan regions. The first pixel seems to have a statistic similar to the rest of the pixels of the overscan. In this case, $\epsilon<5\times 10^{-5}$.
@@
  • Moderate NRI. Figure 8b shows the case where the first pixel in the overscan starts growing as a function of the number of amplifiers proportionally, such as in Eq.(17). In this case, $\epsilon=5\times 10^{-3}$ from the fitted model, which is shown with the black dots.
@@
  • Large NRI. Figure 8c shows an extreme case where the PS and VH low-level voltages of the clocks are set close to the gate voltage of the sense node to force a bad charge removal from the sense node (very large NRI). The intermediate pixels in the transitions from the active to the overscan regions are due to this poor removal capability. In this case, the NRI is very severe, and the pixels in the overscan do not show a proportional growth with the amplifier number. In this extreme case, the best-fit model was obtained by assuming a $\epsilon$ with proportional dependence $\epsilon=2.26\times 10^{-6}Q+3\times 10^{-4}$. The algorithmic model, presented in Appendix B, was used for the calculation.
@@
Figure 8: Measured charge value as a function of the image’s column index for the sixteen MAS-CCD channels. In particular, the plot shows the transition between the pixel in the active region and the overscan region. The different plots show different levels of NRE effect on the measured signal.
@@
Another important aspect to note is that the removal of the charge from the sense nodes requires voltages below the sense node reference voltage. For the case of CCD collecting holes, this means that the voltages should be more negative than the sense node gate voltage. The current limitation of our hardware does not allow us to explore clock voltages below -10 volts; there is an opportunity for future measurements to improve the NRE by using smaller clock values. Another aspect to incorporate in future instrumentation of this kind of sensor is the use of a clock to set voltages in the sense node instead of a fixed bias voltage (as currently being used). A clock would allow an increase in the voltage in the sense node when transferring the change to increase the gradient towards the PS gate.
@@ 6 Conclusions
Code, Data, and Materials Availability
The code to compute the NRE (Sec. 4) is presented in the Appendix B. The data that supports some of the findings of this article are not publicly available due to the large size of datasets and the individual files and privacy concerns. They can be requested from the corresponding author’s email address.
@@ Acknowledgments
The fully depleted Skipper-CCD was developed at Lawrence Berkeley National Laboratory, as were the designs described in this work. The CCD development work was supported in part by the Director, Office of Science, of the U.S. Department of Energy under No. DE-AC02-05CH11231. The multi-amplifier sensing (MAS) CCD was developed as a collaborative endeavor between Lawrence Berkeley National Laboratory and Fermi National Accelerator Laboratory. Funding for the design and fabrication of the MAS device described in this work came from a combination of sources including the DOE Quantum Information Science (QIS) initiative, the DOE Early Career Research Program, and the Laboratory Directed Research and Development Program at Fermi National Accelerator Laboratory under Contract No. DE-AC02-07CH11359.
This research has been partially supported by a Detector Research and Development New Initiatives seed grant at Fermilab, a grant from the Heising-Simons Foundation (#2023-4611), and Javier Tiffenberg’s and Guillermo Fernandez Moroni’s DOE Early Career research programs.
@@ Appendix A Impulse and frequency response of the MAS-CCD to common noise
The Dual-Slope Integrator (DSI) impulse response and frequency response are [16, Chapter 6, pp.574-581][14]
$h_{DSI}\left(t\right)=\begin{cases}\dfrac{A}{T_{s}},\hskip 5.69046ptt_{0}\leq t\leq t_{0}+T_{s}\\ -\dfrac{A}{T_{s}},t_{0}+T_{s}\leq t\leq t_{0}+2T_{s}\\ 0,\hskip 28.45274pt\text{in other case}\end{cases}$ (19)
@@
where $T_{s}$ is the integration time for one sample and $t_{0}$ is an arbitrary start point of the integration. For the frequency response is assumed $t_{0}=0$ without loss of generality.
@@
$H_{SKP}\left(f\right)=\dfrac{H_{DSI}\left(f\right)}{N_{s}}\left|\frac{\sin\left(2\pi N_{s}T_{s}f\right)}{\sin\left(2\pi T_{s}f\right)}\right|$
@@
h_MAS( t) =1Na1Ns∑^N_a-1_m=0∑^N_s-1_n=0h_DSI( t-mΔt -2nT_s) where $\Delta t$ model the required transfer time and intermediate readout time in between stages.
@@
Appendix B NRE recursive algorithm
NRE recursive algorithm that was used to compare with the analytical calculations in Sec. 4. In this complete and fully functional example, the code functions are implemented for the case when the NRE is not a constant. The algorithm also provides a way to implement NRI values which are dependent on the charge in the sense node. The following code produces the measured charge by the amplifiers when a line of the active region is read and 10 overscan pixels, with $N_{a}=16$. It is assumed that the pixels in the active region have 62000e- of charge. The code accounts for 28 pre-scan pixels, before the first amplifier in the chain, and 15 pixels between amplifiers.
Listing 1: NRI implementation in python.
@@ Appendix B NRE recursive algorithm
import matplotlib.pyplot as plt
def pixelsAfterNStages(pix, Na, NaPix):
@@
measPix = np.zeros((Na, Npix + (Na-1)*NaPix))#here we also add
# the pixels in the growing pre-scan for each amplifier
@@
print(n)
print(inPix[n])
iColIndex = j*NaPix #the index accounts for the prescan of
#each amplifier
measPix[j, iColIndex + n] = inPix[n]
# Node Removal Inefficiency calculation (NRI). epsilon is
# the NRI value as used in the text of the article.
@@
#run the software from here
Npre = 28 #the number of pre-scan pixels before the first amplifiers
# in the chain
NpixAct = 256 #the width of the active area of the sensor
NpixOver = 10 #the desired number of pixels in the overscan
Q = 62000 #[e-] amount of charge collected by the pixels in the active
#region
Na = 16 #number of amplifiers in the serial register
NaPix = 15#number of pixels between output stages
pix = np.concatenate((np.zeros(Npre), np.ones(NpixAct)*Q,
np.zeros(NpixOver)), axis=0) #vector of the pixels values in a row of
#the sensor considering prescan of the first amplifier and overscan
print(pix[30])
print(len(pix))
measPix = pixelsAfterNStages(pix, Na, NaPix) #matrix with the pixel
#information of measured by each amplifier
#plot the measured pixels by each amplifier
x = range(np.size(measPix,1))
for i in range(measPix.shape[0]):
plt.plot(x, measPix[i], ’o--’,label=f’Channel{i+1}’,linewidth=1)
plt.xlabel(’Columnindexintheoutputimage’)
plt.ylabel(’Amountofcharge(e-)’)
plt.title(’Pixelvaluesmeasuredbyeachamplifier’)
plt.legend()
plt.show()